OpenStudy (seratul):

Expand the following:

OpenStudy (seratul):

\[\log \sqrt[3]{M^4N}\]

OpenStudy (zrock):

sorry dude still in 8th grade wish i could help

OpenStudy (zrock):

also i have no idea what this means

OpenStudy (seratul):

It's okay. :P

OpenStudy (seratul):

@zepdrix

OpenStudy (eliesaab):

\[ \sqrt[3] {A B}=A^{1/3} B^{1/3} \]

OpenStudy (eliesaab):

\[ \sqrt[3] {A^m b^n}=A^{m/3}B^{n/3} \]

OpenStudy (seratul):

Alright So we can have logM^(4/3)+logN^(1/3)

OpenStudy (seratul):

What would I do after that?

OpenStudy (eliesaab):

Sorry, Let me do it in an easier way \[ \ln(\sqrt[p]{A^kB^s})=\frac 1 p \ln(A^k B^s)=\frac 1 p \left( k \ln(A)+ s \ln(B) \right) \]

OpenStudy (eliesaab):

Use the above formula and you will be done

OpenStudy (eliesaab):

Did you get it?

OpenStudy (seratul):

What about the B^8??

OpenStudy (eliesaab):

It is B^s the letter s

OpenStudy (eliesaab):

In your case \[ \ln \sqrt[3]{M^4N} \] and the general case is \[ \ln(\sqrt[p]{A^kB^s})=\frac 1 p \ln(A^k B^s)=\frac 1 p \left( k \ln(A)+ s \ln(B) \right) \]

OpenStudy (eliesaab):

So p=3, A=M, B=N, k=4 and s=1

OpenStudy (eliesaab):

You should be able to finish it

zepdrix (zepdrix):

Oh log expansion :) These are fun

zepdrix (zepdrix):

You figure it out?